If you could help me with this it will be greatfully appreciated and i will donate to your site as soon as possible.
i tryed to do these problems over and over and i really just dont get it.
its really confusing thank you.
Kacey

1. find 60% of 150?
2. what number is 40% of 95?
3. 21 is 70% of what number?
4. find 20% of 120?
5. find 7% of 80?
6. 63 is 60% of what number?
7. 12 is 30% of what number?
8. 90 is 45% of what number?
9. what number is 27% of 50?
10. what number is 70% of 122?
11. what number is 12% of 85?
12. find 14% of 150?
13. 26 is 65% of what number?
14. what number is 67% of 140?
15. 108 is 90% of what number?
16. fidn 34% of 85?
17. 50 is 25% of what number?
18. what number is 95% of 90?
19. 21 is 35% of what number?
20. find 22% of 55?
21. find 14.5% of 500?
22. 4 is 0.8% of what number?

1. 72 is what percent of 90?
2. find 10% of 240?
3. 90 is what percent of 360?
4. find 8% of 120?

Answer by checkley75(3666)

1) .60*150=90
2) .40*95=38
3) .21*70=14.7
4) .20*120=24
HERE'S A SAMPLE OF THE PROCEDURE.
NOW TRY TO SOLVE THE REST.